Scrooge and Zilch, a public accounting firm in London, is engaged in the preparation of income tax returns for individuals. The firm uses the weighted-average method of process costing for internal reporting. The following information pertains to February. (£ denotes the British monetary unit, pounds sterling.)*

Returns in process, February 1:
(35% complete) 450
Returns started in February 1,200
Returns in process, February 28:
(60% complete) 490
Returns in process, February 1:
Labor £ 3,900
Overhead 5,000
Labor, February (9,853 hours) 118,236
Overhead, February 56,068

*Although the euro is used in most European markets, day-to-day business in the United Kingdom continues to be conducted in pounds sterling.

2. Compute the cost of returns in process as of February 28.
Answer #1

Unit transferred out = 450+1200-490 = 1160

Equivalent unit = 1160+(490*60%) = 1454

Cost per equivalent unit = (3900+5000+118236+56068)/1454 = 126

2) Cost of returns in process = 490*60%*126 = $37044